Run Planning Meeting Notes, April 24 - April 30, 2025
From GlueXWiki
↑ Back to Run Coordination Meetings:Spring2025#Runplan
Contents
Tuesday April 29
Shift Crew Summaries
Reports
- ROCTAGH1 power cycled; DAQ live again.
- 13:00; ion chamber calibration on diamond
- JD80-212 Diamond preparation/setup for production New setup
- 15:18: BCAL/FCAL LED pulsing restarted
- Trigger
- Detector Experts - please examine timing windows/cuts
- Trigger Energy and follow-up
- CDC
- BCAL
- Added to production trigger; monitoring plot updates required.
- Check of BCAL-TAGH/TAGM timing
- FCAL
- TOF
- TRD
Issues
- TRD removal date: Thu.
- Incorrect fan tray used in swap yesterday; correcting today.
- Gas bottle swaps in hall (He on GC, Ar)
- Calibration factor adjustment on Tagger e- beam dump H2O conductivity.
- All detectors: Check timing cuts/windows in preparation for production
- hdmon occupancy stalls w/o reason
Plans
- Upcoming Plans:
- Full survey - Hall to Restricted access as soon as is practicable (Accommodating gas bottle swaps etc.)
- Full survey - Tagger to Controlled access (Beam dump H2O conductivity cal. factor adjustment)
Monday April 28
Shift Crew Summaries
Reports
- IBCAD00C temp alarm; AES-IC replaced controller. Dipped again; eyeing relay box swap.
- Day shift access (10:30-12:00) for several repairs
- Trigger development (16:00-19:15)
- Beam will cut off at 7:15 on Wed. for maintenance day. Projected(8 hrs): AM: Wien angle change/RF recovery; PM: beam studies
- ECAL
- HV board trip 3rd board on d1-4-top crate (board 2 counting from 0); and again
- Successful remote power reset of D1-4-TOP following power reset pulse signal modification. Changed for remaining crates.
- HV module in slot 2 of the D1-1-TOP crate swapped.
- FCAL
- FCAL gains following last round of HV adjustments.
- Next round of HVs loaded after BCAL in trigger and overview after power cycle.
- Trigger Development
- FDC
- Wire efficiency status
- Drift times for cathode and wires
- Access found MFC calibrated for Argon; not Nitrogen. Gas factor adjusted. Fill cycle 1392s
- Low RPMs FDC5 fan tray
- TOF
- TRD
Issues
- Trigger / PS rates not in EPICS from 2:47 am; perhaps related to data rate loss at 2:43
- 2 MeV energy oscillation starting 6:00 4/28.
- Kr gas supply - ~4 days remaining.
- Gas chromotography He bottle empty
Plans
- Upcoming Plans:
- Diamond preparations
- CDC preparations
Weekend April 26-April 28 (Owl)
Shift Crew Summaries
- April 26: Owl (0000-0800) Day (0800-1600) Swing (1600-2400)
- April 27: Owl (0000-0800) Day (0800-1600) Swing (1600-2400)
- April 28: Owl (0000-0800)
Reports
Saturday
- 10 a.m. RC Meeting
- Access to reset ZDC prototype DAQ computers
- 19:00 Ecal major alarms; 19:48 Rapid access by Sasha to troubleshoot. Beam back 20:43.
- 23:00: Tagger NMR softioc reset for comm. failure - OK.
- FDC
- FDC pressure during gas sampling
- Gas Analysis Summary and follow-up; points to bad CO2 regulation.
- Switched I/O channel on CO2 channel - no dice; suggests bad mass-flow controller. Present gas mix calculated and adjusted in interim.
- Detector voltage lowered to prevent HV trips with new gas mix. Evolving gas levels monitored on FDC-1 with Gas Chromatograph.
- Perhaps replace both MFCs
- FCAL
- New voltages starting with run 131217
- Tracks extrapolated into FCAL
- ECAL
- HV drop - Run #131226
- Two ECAL modules 'externally disabled'/locked up: D1-4-TOP, Slot 2 & D1-4-TOP, Slot 4. Rapid access to power cycle crate recovered.
- Voltage GUI during alarms
- TRD
Sunday / Mon Owl
- DAQ issues starting ~2:30am.
- Rootspy fatal errors
- https://logbooks.jlab.org/entry/4366230 Tagger vacuum issues starting ~7:30] HD mechanical called in. AD00A recovered; AD00B/C rising.
- Al target Ion chamber calibration
- Arrived & examined AD00B pump (downstream of Tagger) - speed dropping. Access & controller reset; area pumped down again.
- Trigger some difficulty locating main bunch - timing RF-to-ECAL effect? 3 innermost ECAL layers excluded.
- FDC
- FDC-5 unresponsive; rapid access to restore communication. New DAQ setup file w/o FDC5 (unresponsive) ~3:30 am, hall access needed/authorized to restore FDC-5 communication. DAQ expert arrived 4:50; found FDC5 crate fan failure - crate does not stay powered on more than 30 seconds. replacement of the fan tray on the crate is needed to bring the crate back up. Will run without it until the fan tray can be replaced
- Sun. day gas chromotography
- FCAL
- Base (16,-6) locked up. Reset procedure performed. Voltages came back differently - restore procedure in wiki modified to correct. Production HV module loaded.
- New voltages loaded ~15:00 - run 131237 onwards.
- ECAL
- Rapid Access for HV crate local reset on day shift Remote power reset of the crates remains an issue - delay adjustment required?
- [https://logbooks.jlab.org/entry/4366356 ROCECAL4 FADC Slot 20 (D1-2-TOP) replaced ACDI-481 with ACDI-307 for readout issues.
- ECAL HV down for channels 20,12 through 20,9 to 20,9 through 9,9. Unable to remote reset.
- PS
- TRD
Issues
Saturday
- ECAL HVs unable to remote reset.
- FDC input change unhelpful - gas mix change adjusted. MFC valves perhaps both require swap.
Sunday
- Crate fan tray supporting FDC-5 failed; repair required. Bypassed in trigger for now.
- VTPAD00B local controller reset to recover from error.
- ECAL HVs unable to remote reset.
- FADC swap in ROCECAL4.
- FCAL reloads perhaps not restoring HV properly - monitor and use file if necessary
Monday Owl
- ECAL HVs unable to remote reset.
Plans
- Trigger commissioning
Friday April 25
Shift Crew Summaries
Reports
- Some beam current instabilities across Runs #131201 & 131202 on day shift; perhaps due to INJ magnet change. Reverted by Ops.
- 15:45-17:00 Rapid Access - snoop detector gas concentrations.
- Some errors observed on owl shift when attempting to start a new run.
- Hydra issues at end of owl shift - monitoring process restarted
- Trigger:
- Testing on swing shift
- ECAL & FCAL trigger: Bit 0
- BCAL trigger (prescaled): Bit 2
- PS trigger: Bit 3
- Switch to ECAL/FCAL trigger 20:08 First run: 131207
- FDC:
- Drift time comparisons
- Gain change and Follow-up
- Effects of +50V change: rho yield x5; phis x4; & omegas x2.5. I-beam oscillating 240-270nA < 300 requested.
- Follow-up: ~15% efficiency gain, but still << 2023.
- Pulse height v time comparison
- Time-to-distance comparison and follow-up
- Gas Mixture Summary log entry
- FDC-1 Gas mixture comparison - 2021: 30% Ar/70%CO2; now 22%/78%. CDC should be 50/50; measure both overnight to calibrate analyzer.
- Next measurement on FDC-1: 22.6%/77.4% - No traces of N2.
- Similar 3rd FDC-1 measurement. Switching gas analyzer input over to CDC at 19:17
- FDC-2 - 22.6% Ar /77.4% CO2
- Track Match to FCAL and TOF run 131206 (w/ FCAL only Trigger)
- ECAL:
- HV module issue. D1-4-MID crate; Slot 12 - not visible by GECO; remote IOC reset/crate power cycle unsuccessful. Manual power cycle under access recovered.
- HV errors; reboot and re-ramp to clear.
- FCAL:
- Omega > 3pi peak
- FCAL base lockup for issue channel (-17,11)
- NEW RC values for CAEN TDCs in config file
- Base lock-up on 3 channels: (15,-10)/(-1,-11)/(-10,-11) lost communication; reset at end of run #131213.
- TDC:
- CDC:
- Alcohol bubbler topped off
- Pedestals after yesterday's baseline recalibrations.
- TOF:
- Occupancy follow-up - Model retraining perhaps necessary.
Issues
Plans
- Upcoming Plans:
- Parasitic setup of MPGD Install complete - ready to take data.
Thursday April 24
Shift Crew Summaries
Reports
- Beam back to hall 9:50 am 4/24 following Tagger PS repair.
- Accelerator Fast Feedback study to Hall A from 9-11 a.m.
- ~11 a.m.: Left CO2 gas bank at 300 PSI - ready for switch over
- Solenoid
- FDC:
- ECAL
- Voltages alarm at 19:53; Sasha rapid access - power cycle crate for flakey HV module; fixed. DAQ restarted & FCAL voltages restored. (Runs 131193 & 131194 junk.) Return to production 22:15.
- ECAL voltages adjusted; iteration 5 Pi0 peak is now clearly visible in ALL ECAL channels, except those obstructed by an absorber.
- Occupancy Col = -13 Row = 3; suspicious module.
- ECAL/BCAL distribution tail - Asymmetric Timing plots
- CDC
- PS
- FADC Scaler Readouts: Left Arm, scaler 2 & 28 off. Right Arm, scaler 4 is off.
- TAGH
- Scaler readouts: FADC, Disc Readout1 and Disc Trigger1, scalers 21 and 50 are off.
Issues
- Many 2L22 HV trips in CEBAF overnight
- CEBAF drops from beam to power permit several times overnight due to loss of A-chain permit on beam current monitor loss integration units.
- Run 131183:
- FCAL base (5:-21) lockup
- Gas shed upper alcohol bubbler ready for refill.
Plans
- Upcoming Plans:
Wednesday April 23
Shift Crew Summaries
Reports
- Maintenance day:
- SAF113 training (8+1 person)
- Radcon survey for Restricted Access
- FOPT before maintenance day to Hall D
- CEBAF Maintenance
- Calorimeter Work
- FCAL chiller LCW topped off; crane inspection complete
- BCAL/FCAL LED pulsing script disabled - pending FCAL commissioning completion
- 4 FCAL bases changed; databases updated
- FCAL HV updated/loaded; new FCAL/BCAL pedestals
- ECAL/FCAL Time & Energy calibrations
- FCAL new HV plot
- FCAL channel 13:-3 locked for run 131181
- FDC discriminator threshold lowered to 8V
- CDC anomalous region in Cosmic run 131169
- ECAL voltages adjusted: iteration 4 - 170 modules for Pi0 peak visibility
- Run 131178: New missing ECAL channnel - Low rate in (-5,-5)
- Harp Scans:
- Return to production ~1:40 a.m. 4/24 following harp scans
- Parasitic work / experiments:
- EIC + MOLLER (how many people?)
- Open the Hall gate
- Disassemble parts, Crane use
- SVT Test Stand/Telescope Removal
- Change Management Video production (Doug H.++)
- Wrap up by pm
- Parasitic work / experiments:
Issues
- Beam Position:
- Beam Energy
- Solenoid status:
- BCAL Occupancy plot:
- Tagger drop (~6:50 am):
- Many 2L22 HV trips in CEBAF overnight
Plans
- Upcoming Plans:
Tuesday April 22
Shift Crew Summaries
- Day 0800-1600 (Nice template developed by Kevin Scheuer)
- Evening 1600-2400
- Owl April 23, 0000-0800
Reports
- ECAL missing \pi^0 problem: solved?
- ECAL HV adjusted HallD Run #131151: Ecal scaler rates
- Updated a few rootspy references with run 131153
- new RootSpy plot added: ECAL-RF timing
Issues
Plans
- Maintenance day:
- SAF113 training (8+1 person)
- Radcon survey for Restricted Access
- Parasitic experiments
- EIC + MOLLER (how many people?)
- open the Hall gate
- Dissasemble parts, Crane use
- need to wrap up by 3-4pm
Previous RC
On-going Plans
- Short term run plan
- highest priority: ECAL+FCAL+BCAL trigger preparing for pi0 production
- Diamond alignment
- Production: pi0 running
- anything else?